OpenStudy (anonymous):

can you explain 4.5(0.2)

OpenStudy (tkhunny):

What is to be done? Multiply? 4.5*0.2 = \(\dfrac{9}{2}\cdot\dfrac{1}{5} = \dfrac{9}{10} = 0.9\)

OpenStudy (anonymous):

Thank You